OpenStudy (anonymous):

The function f(x) = 4(3)^x represents the growth of a dragonfly population every year in a remote swamp. Erin wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for Erin's purpose, and what is the new growth rate? f(x) = 4(3)x; growth rate 300% f(x) = 4(3)x, growth rate 4% f(x) = 4(1.32)x; growth rate 4% f(x) = 4(1.32)4x; growth rate 32%

OpenStudy (anonymous):

4 (3)^x = 4(3)^(4x/4) = 4 (3^(1/4))^(4x) = 4 (1.316)^(4x) rate = 31.6% (or about 32% according the choices)

OpenStudy (anonymous):

thank youu!!